A rhinoceros is at the origin of coordinates at time t1=0. For the time interval from t1 to t2 = 13.0 s , the rhino's average velocity has x-component -3.2 m/s and y-component 6.0 m/s .

At time t2= 13.0 s what is the x-coordinate of the rhino?

At time t2= 13.0 s what is the y-coordinate of the rhino?

How far is the rhino from the origin?
Answer #1

2 dioection distance = -3.2 X13 = -4.6m g direction distance - 6x13 = 78m Distance = √(x² + 162 Distance = V6-41633+ (7872 =